# Sparse Extended Mean-Variance-CVaR Portfolios with Short-selling

### v1

@@ 4 Numerical Results
In this section, we compare the performance of Algorithm 2 with the algorithm in [7]. The direct approach of solving ($P$) by CVX-Mosek is also reported for gap computing. We use the data of S&P index for 2018-2021 with 120 stocks when $\lambda_{1}=\lambda_{2}={1}/{3}$, $\phi=0$, $U_{i}=-L_{i}=0.2;\forall i\in[n],\delta=0.002$ and $\beta=0.95$. Moreover, we apply the Geometric Brownian Motion (GBM) model in [1] to generate scenarios. In the PADM algorithm of [7], the initial penalty parameter is set to $1.2$ and it is updated with the factor 3. The inner loop is stopped when $||(x^{s,l},w^{s,l})-(x^{s,l-1},w^{s,l-1})||_{1}\leq 10^{-5}$. It terminates with a partial minimum if $||x-w||_{1}\leq 10^{-5}$. In Algorithm 2, the initial penalty parameter is set to $1.2$ as well and it is updated with the factor 3. In this Algorithm, the inner loop is stopped when $\epsilon_{I}=10^{-5}$ and the outer loop is stopped when a convergence criterion is met $\epsilon_{O}=10^{-5}$. All computations are performed in MATLAB R2017a on a 2.50 GHz laptop with 4 GB of RAM, and CVX 2.2 ([5]) is used to solve the optimization models. The results comparing returns, risks, Sharpe ratios, CVaR values, CPU times, and gaps for different $m$ values are reported in Table 1. The gap in this table is ${|f-\overline{f}|}/{(|f|+1)}$ where $f$ is the return (risk, Sharpe ratio, CVaR) for direct solution approach and $\overline{f}$ is the return (risk, Sharpe ratio, CVaR) for the PADM algorithm or Algorithm 2.
These results show that both Algorithm 2 and PADM are extremely faster than the direct solution approach. Also, Algorithm 2 is about twice as fast as the PADM while having competitive gaps of returns, risks, CVaR, and Sharpe ratios; for all $m$ values. These results for $m=3000$ also are depicted in Figure 1 for different numbers of stocks showing the competitiveness of Algorithm 2. These results confirm that Algorithm 2 is a better alternative to the direct solution approach than the PADM in [7].
@@ 5 Conclusion
In conclusion, this paper designs a penalty decomposition algorithm customized for tackling the challenging sparse extended mean-variance-CVaR portfolio optimization problem in finance. This algorithm needs to solve a sequence of penalty subproblems. By employing a block coordinate method, we adeptly exploit every structure in the problem to manage each penalty subproblem, establishing their well-posed nature and deriving closed-form solutions wherever feasible. Our comprehensive convergence analysis demonstrates the efficacy of our introduced algorithm in efficiently reaching a local minimizer of this non-differentiable and nonconvex optimization problem. Furthermore, extensive numerical experiments conducted on real-world datasets validate the practical applicability, effectiveness, and robustness of our introduced algorithm across various evaluation criteria. Overall, this research contributes significantly to the field of portfolio optimization by offering a novel and efficient solution approach with promising practical implications in finance.
Table 1: Comparison of returns, risks, Sharp ratios, CVaR and CPU times for the direct approach, PADM, and Algorithm 2 for different $m$ values for S&P index data with different confidence levels $\beta$ and different number of scenarios $m$ when $\lambda_{1}=\lambda_{2}=\dfrac{1}{3}$, and $k=30$.
@@
Figure 1: Comparison of returns, risks, Sharpe ratios, and CVaR for different $k$ values for 130 stocks of S&P index.

### v2

@@ 4 Numerical Results
In this section, we compare the performance of Algorithm 2 with the algorithm in [9]. The direct approach of solving ($P$) using CVX-Mosek is also reported for gap computation. We use the data of the S&P index for 2018-2021, with 120 stocks when $\lambda_{1}=\lambda_{2}={1}/{3}$, $\phi=0$, $U_{i}=-L_{i}=0.2;\forall i\in[n],\delta=0.002$, and $\beta=0.95$. Moreover, we apply the Geometric Brownian Motion (GBM) model in [1] to generate scenarios. In the PADM algorithm of [9], the initial penalty parameter is set to $1.2$, and it is updated by a factor of 3. The inner loop is stopped when $(x^{s,l},w^{s,l})-(x^{s,l-1},w^{s,l-1})||_{1}\leq 10^{-5}$. It terminates with a partial minimum if $x-w||_{1}\leq 10^{-5}$. In Algorithm 2, the initial penalty parameter $\rho$ is set to $1.2$, and it is updated with the factor $r=3$. The inner loop is stopped for $\epsilon_{I}=10^{-5}$, and the outer loop is terminated when the convergence criterion $\epsilon_{O}=10^{-5}$ is met. All computations are performed in MATLAB R2017a on a 2.50 GHz laptop with 4 GB of RAM, and CVX 2.2 ([8]) is used to solve the optimization models. The results comparing returns, risks, Sharpe ratios, CVaR values, CPU times, and gaps for different $m$ values are reported in Table 1. The gap in this table is ${|f-\overline{f}|}/{(|f|+1)}$, where $f$ is the return (risk, Sharpe ratio, CVaR) for the direct solution approach, and $\overline{f}$ is the return (risk, Sharpe ratio, CVaR) for the PADM algorithm or Algorithm 2.
These results show that both Algorithm 2 and PADM are significantly faster than the direct solution approach. Further, Algorithm 2 is about twice as fast as the PADM while having competitive gaps in returns, risks, CVaR, and Sharpe ratios for all $m$ values. These results for $m=3000$ are also depicted in Figure 1 for different numbers of stocks, demonstrating the competitiveness of Algorithm 2. These results confirm that Algorithm 2 is a better alternative to the direct solution approach than the PADM in [9].
@@ 5 Conclusion
In conclusion, this paper designs a penalty decomposition algorithm customized for tackling the challenging sparse extended mean-variance-CVaR portfolio optimization problem in finance. This algorithm needs to solve a sequence of penalty subproblems. By employing a block coordinate method, we adeptly exploit every structure in the problem to manage each penalty subproblem, establishing their well-posed nature and deriving closed-form solutions wherever feasible. Our comprehensive convergence analysis demonstrates the efficacy of our introduced algorithm in efficiently reaching a Lu–Zhang minimizer of this non-differentiable and nonconvex optimization problem. Furthermore, extensive numerical experiments conducted on real-world datasets validate the practical applicability, effectiveness, and robustness of the algorithm we introduced across various evaluation criteria. Overall, this research contributes significantly to the field of portfolio optimization by offering a novel and efficient solution approach with promising practical implications in finance.
Figure 1: Comparison of returns, risks, Sharpe ratios, and CVaR for different $k$ values for 130 stocks of the S&P index.
Table 1: Comparison of returns, risks, Sharpe ratios, CVaR and CPU times for the direct approach, PADM, and Algorithm 2 for different $m$ values for S&P index data with different confidence levels $\beta$ and different number of scenarios $m$ when $\lambda_{1}=\lambda_{2}=\dfrac{1}{3}$, and $k=30$.